OpenStudy (anonymous):

Suppose the volume of milk in a glass is 230 cubic cm. If 1 gallon of milk equals 3.79 liters, how many glasses of milk are in 1 gallon

OpenStudy (anonymous):

So 1mL=1cm^3 so 230cm^3=230mL and 1000mL = 1L so 230mL=0.23L 1gallon =3.79L so \[1 gal *\frac{ 3.79L }{ 1gal }*\frac{ 1000mL }{ 1L }*\frac{ 1cm ^{3} }{ 1mL }*\frac{ 1glass }{ 230cm ^{3} }=16.478glasses\]

OpenStudy (anonymous):

You would round down to get 16 glasses